OpenStudy (anonymous):

log4 (2x+3) - 2log4 x= 2 4 is base

OpenStudy (anonymous):

answer is 9

OpenStudy (anonymous):

what i dont understand

OpenStudy (anonymous):

log_b(A) - log_b(B) = log_b(A/B) log_b(A) + log_b(B) = log_b(AB) log_b(A^B) = B*log_b(A)

OpenStudy (anonymous):

why nine 9

OpenStudy (anonymous):

\[\log_4(2x + 3) - \log(x^2) = 2\]\[\log_4\left( \frac{ 2x+3 }{ x^2 } \right) = 2\]\[\frac{ 2x+3 }{ x^2 } = 4^2\]\[16x^2 - 2x - 3 = 0\]solve for x

OpenStudy (anonymous):

he's lying. he thinks he's funny

OpenStudy (anonymous):

ah ok .. thank you

OpenStudy (anonymous):

i think i will use quadratic equation

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

good idea. i would too

OpenStudy (anonymous):

can i use quadratic equation ?

OpenStudy (anonymous):

with a second order polynomial: always

OpenStudy (anonymous):

where did you get the 16x^2

OpenStudy (anonymous):